a force 2i+5j displaces a body from (2,1,1) to (1,2,3). find the work done if the force is in newtons and displacement is in meters?

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Force F = (2i + 5j)
d S = (1,2,3) - (2,1,1) = - i + j + 2k
Now work done = F. dS = -2 +5 = 3 units
Regards
Arun (askIITians forum expert)